# The norms for symmetric and antisymmetric tensor products of the weighted shift operators

### v1

@@ 3.2 Proof of “$(1)\Leftrightarrow(3)$
$\Sigma_{n,\mathfrak{i}}^{\prime\prime\prime}=\left(\mathop{\bigsqcup}_{t=1}^{\mathscr{M}}\mathop{\bigcup}\limits_{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},t}\backslash\mathcal{E}_{k+S_{\mathfrak{l}},t}^{\prime}}\mathcal{N}_{\mathfrak{j},\mathfrak{i}}\right).$ (3.51)
@@
$\left\|S_{{l_{1}}}\wedge S_{{l_{2}}}\wedge\cdots\wedge S_{{l_{n}}}|_{H_{d-1}\odot\cdots\odot H_{d-1}}\right\|\geq 1.$ (3.52)
For $d\leq i_{1}<\cdots<i_{n}$ with $|\mathfrak{i}|=k$, let $a_{\mathfrak{i}}=\frac{1}{\sqrt{W_{(k,n)}}}$. Obviously $\sum\limits\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}$ is a unit vector. To calculate the norm of $S_{{l_{1}}}\wedge\cdots\wedge S_{{l_{n}}}$, we will estimate the norms of
$S_{{l_{1}}}\wedge\cdots\wedge S_{{l_{n}}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right).$
@@
$\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}=0~{}\text{or}~{}e_{i_{1}+l_{\pi(1)}}\wedge\cdots\wedge e_{i_{n}+l_{\pi(n)}}=0.$
@@
$\left(S_{{l_{1}}}\wedge\cdots\wedge S_{,{l_{n}}}\right)\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}\frac{e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}}{\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}$ (3.53)
@@
$~{}+\sum\limits_{t=1}^{\mathscr{M}}\sum\limits_{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}$
@@
$~{}+\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}$
$~{}+\sum\limits_{\mathfrak{j}\in\check{\mathcal{A}}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}$
@@
$\left\|\left(S_{{l_{1}}}\wedge\cdots\wedge S_{{l_{n}}}\right)\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}\frac{e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}}{\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right\|^{2}$ (3.55)
$\geq$ $\left\|\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}\right\|^{2}$
$=$ $\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)^{2}\left\|e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}\right\|^{2}$
$=$ $\frac{1}{n!}\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)^{2}$
$\geq$ $\frac{{}^{\#}\tilde{\mathcal{A}}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}^{\prime}}{W_{(k,n)}}\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)},$
where the last inequality comes from (3.32) and Lemma 3.7, together with the fact that $\mathfrak{i}\in\mathcal{R}_{\mathfrak{j}}$ if and only if $\mathcal{N}_{\mathfrak{j},\mathfrak{i}}\neq\emptyset.$ Thus, by Lemma 3.6, (3.55), and the arbitrariness of $\varepsilon$, we conclude that (3.52) holds. $\square$
@@ 4 Lower bound of the norm for symmetric tensor product
This section is devoted to solving [6, Problem 1 and Problem 2]. At first, recall that [6, Problem 1 and Problem 2] are stated as follows.
@@
([6, Problem 1]) For $x_{1},x_{2},\cdots x_{n}\in\mathcal{H}$, is
@@
([6, Problem 2]) For $\mathbb{A}_{1},\mathbb{A}_{2},\cdots,\mathbb{A}_{n}\in\mathcal{B(H)}$, is
@@
$\left\|x_{1}\odot x_{2}\odot\cdots\odot x_{n}\right\|^{2}\geq\frac{1}{n!}\prod\limits_{j=1}^{n}\sum\limits_{i_{j}=0}^{\infty}\left|a_{j,i_{j}}\right|^{2}=\frac{1}{n!}\prod\limits_{j=1}^{n}\left\|x_{j}\right\|^{2}.$
Proposition 4.4.
@@
$\geq\frac{1}{\sqrt{n!}}\sup_{x\in\mathcal{H}\atop\left\|x\right\|=1}\left\{\left\|\mathbb{A}_{1}x\right\|\left\|\mathbb{A}_{2}x\right\|\cdots\left\|\mathbb{A}_{n}x\right\|\right\}.$
Acknowledgement. This work is supported by National Natural Science Foundation of China (No. 11871308, 12271298).

### v2

@@ 3.2 Proof of $(1)\Leftrightarrow(3)$
[…]
@@
$\left\|S_{{\alpha,l_{1}}}\wedge S_{{\alpha},l_{2}}\wedge\cdots\wedge S_{{\alpha,l_{n}}}|_{H_{d-1}\odot\cdots\odot H_{d-1}}\right\|\geq 1.$ (3.52)
For $d\leq i_{1}<\cdots<i_{n}$ with $|\mathfrak{i}|=k$, let $a_{\mathfrak{i}}=\frac{1}{\sqrt{W_{(k,n)}}}$. Obviously $\sum\limits\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}\frac{e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}}{\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}$ is a unit vector. To calculate the norm of $S_{{\alpha,l_{1}}}\wedge S_{{\alpha,l_{2}}}\wedge\cdots\wedge S_{{\alpha,l_{n}}}|_{H_{d-1}\odot\cdots\odot H_{d-1}}$, we will estimate the norms of
$S_{{\alpha,l_{1}}}\wedge S_{{\alpha,l_{2}}}\wedge\cdots\wedge S_{{\alpha,l_{n}}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}\frac{e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}}{\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right).$
@@
$e_{i_{1}+l_{\pi(1)}}\wedge\cdots\wedge e_{i_{n}+l_{\pi(n)}}=0.$
@@
$\left(S_{{\alpha,l_{1}}}\wedge S_{{\alpha,l_{2}}}\wedge\cdots\wedge S_{{\alpha,l_{n}}}\right)\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}\frac{e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}}{\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}$ (3.53)
@@
$~+\sum\limits_{t=1}^{\mathscr{M}}\sum\limits_{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},t}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}$
@@
$~+\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}$
$~+\sum\limits_{\mathfrak{j}\in\check{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}$
@@
$\left\|\left(S_{{\alpha,l_{1}}}\wedge S_{{\alpha,l_{2}}}\wedge\cdots\wedge S_{{\alpha,l_{n}}}\right)\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}a_{\mathfrak{i}}\frac{e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}}{\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right\|^{2}$ (3.55)
$\geq$ $\left\|\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}\right\|^{2}$
$=$ $\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)^{2}\left\|e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}\right\|^{2}$
$=$ $\frac{1}{n!}\sum\limits_{\mathfrak{j}\in\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}\left(\sum\limits_{d\leq i_{1}<\cdots<i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in\mathcal{N}_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\mathfrak{i}}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{i_{1}}\wedge\cdots\wedge e_{i_{n}}\right\|}\right)^{2}$
$\geq$ $\frac{{}^{\#}\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}{W_{(k,n)}}\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)},$
where the last inequality comes from (3.32) and Lemma 3.7, together with the fact that $\mathfrak{i}\in\mathcal{R}_{\mathfrak{j}}$ if and only if $\mathcal{N}_{\mathfrak{j},\mathfrak{i}}\neq\emptyset.$ Thus, by Lemma 3.6, (3.55), and the arbitrariness of $\varepsilon$, we conclude that (3.52) holds. ∎
@@ 4 Lower bound of the norm for symmetric tensor product
This section is devoted to solving [7, Problem 1 and Problem 2]. At first, recall that [7, Problem 1 and Problem 2] are stated as follows.
@@
([7, Problem 1]) For $x_{1},x_{2},\cdots x_{n}\in\mathcal{H}$, is
@@
([7, Problem 2]) For $\mathbb{A}_{1},\mathbb{A}_{2},\cdots,\mathbb{A}_{n}\in\mathcal{B(H)}$, is
@@
$\left\|x_{1}\odot x_{2}\odot\cdots\odot x_{n}\right\|^{2}\geq\frac{1}{n!}\prod\limits_{j=1}^{n}\sum\limits_{i_{j}=0}^{\infty}\left|a_{j,i_{j}}\right|^{2}=\frac{1}{n!}\prod\limits_{j=1}^{n}\left\|x_{j}\right\|^{2}.\qed$
Proposition $4.$
@@
$\geq\frac{1}{\sqrt{n!}}\sup_{x\in\mathcal{H}\atop\left\|x\right\|=1}\left\{\left\|\mathbb{A}_{1}x\right\|\left\|\mathbb{A}_{2}x\right\|\cdots\left\|\mathbb{A}_{n}x\right\|\right\}.\qed$
Acknowledgement. This work is supported by National Science Foundation of China (No. 11871308, 12271298) and the Taishan Scholars Project (TSTP20250702). The authors thank the referee for helpful suggestions, which make this paper more readable.